# Masked Diffusion Models as Energy Minimization

### v1

@@ Appendix B An Intuitive Explanation of Geodesics
This appendix provides an intuitive introduction to geodesics and exponential maps to clarify the geometric interpretation of MDM in [42]. Readers seeking formal mathematical definitions of these differential geometry concepts may refer to [5] for complete technical specifications.
@@
Recent theoretical advances [42] reveal that MDM’s conditional probability flow in Eq. (3) forms exactly the geodesic curve in spherical geometry (see Fig. 6). The interpolation schedule $\gamma_{t}$ is uniquely determined by the mask schedule $\alpha_{t}$ as $\gamma_{t}=\frac{2}{\pi}\arcsin\sqrt{\alpha_{t}}$, which is equivalent with Condition 3.4.
@@ C.2 Proof of the Invariance of Training Loss to the Mask Schedule
Different equivalence expressions of the training loss of MDM has been proved invariant to the choice of $\alpha_{t}$ in multiple works [12, 38, 36]. We adapt a proof from [36] by examining the negative evidence lower bound (NELBO) through token-level denoising components:
@@
Proposition C.2 (Schedule Invariance, Proof Adapted from [36]).
@@ C.4 Decomposition of Conditional Kinetic Energy along Sequence Dimension
The pivotal decomposition in $(1)$ leverages two structural properties: (i) The conditional probability flow (Eq. (3)) exhibits token-wise independence, and (ii) The conditional rate matrix (Lemma C.3) nullifies transitions altering multiple tokens simultaneously. These enable reduction of full-sequence transitions to single-token operations, with remaining n1n-1 tokens contributing constant combinatorial factors. This structural property persists across various DPF implementations including MDM and Discrete Flow Matching [32, 37], validating the conditional kinetic energy as theoretically sound surrogate objective. Notably, standard kinetic energy $\mathcal{E}$ lacks such decomposition due to $p_{t}(z)$’s dependence on cross-token correlations.
@@ Appendix D Proof of Main Results
Our argument leverages a foundational decomposition from [44] regarding concrete score representations:
@@
where $(1)$ leverages orthogonality $p_{0}\cdot p_{1}=0$. Therefore, we proved that MDM with schedule schedule $\alpha_{t}^{\star}$ generates the minimal-length curve as well as minimal-energy conditional probability path, validating and generalizing the conclusion in [42] from a energy perspective.
@@ D.5 Optimality in the Discretized Case
where $Q_{t|1}(\alpha_{t}^{\star})$ refer to conditional rate matrix in MDM case derived in Appendix C.3 under Condition 3.4. We adapt the following key result from prior analysis  [37]of the $\gamma_{t}=t$ case:
@@ E.2 Standard Benchmarks and Evaluation Settings
Building upon established practices in LLM evaluation [30, 41, 43], we evaluate performance across key dimensions including: general ability (BBH [28]), mathematics (GSM8K [10], Hendrycks MATH [11], Minerva MATH [18]), and code generation (MBPP [7], HumanEval [9]). Evaluation follows the conditional generation paradigm, where models produce completions given task prompts, with performance quantified through exact match or other domain-specific evaluation metrics.
Our implementation leverages the open-source pretrained weights and evaluation toolkit from LLaDA [43], modifying only the mask schedule that governs the iterative unmasking process. The mask schedule affects the number of tokens unmasked at each step, with certain schedules permitting zero-token unmasking during the process. Therefore, generation quality discrepancies occur even when sampling steps are set as the sequence length. All experiments can be efficiently conducted on a single NVIDIA A800 GPU.
@@ E.3 Additional Results and Raw Data
Fig. 7 shows the result of our main experiments on benchmark BBH [28] and GSM8K [10], where our beta-parameterized schedules exhibit comparable yet not better performance than the linear schedule.
@@
Figure 7: Performance evaluation of energy-optimized schedules on BBH [28] and GSM8K [10] where our beta-parameterized schedules exhibit comparable yet not better performance than the linear schedule. Each panel corresponds to a distinct benchmark. The x-axis displays sampling steps on a logarithmic scale, while the y-axis quantifies task performance, where higher values denote superior generation quality.
Table 2: Performance evaluation of beta-parameterized schedules on MBPP [7] benchmark. All experiments fix generation length at 256 and higher values indicate better sampling quality.
@@
Table 3: Performance evaluation of beta-parameterized schedules on HumanEval [9] benchmark. All experiments fix generation length at 256 and higher values indicate better sampling quality.
@@
Table 4: Performance evaluation of beta-parameterized schedules on BBH [28] benchmark. All experiments fix generation length at 128 and higher values indicate better sampling quality.
@@
Table 5: Performance evaluation of beta-parameterized schedules on GSM8K [10] benchmark. All experiments fix generation length at 128 and higher values indicate better sampling quality.
@@
Table 6: Performance evaluation of beta-parameterized schedules on Hendrycks Math [11] benchmark. All experiments fix generation length at 256 and higher values indicate better sampling quality.
@@
Table 7: Performance evaluation of beta-parameterized schedules on Minerva Math [18] benchmark. All experiments fix generation length at 256 and higher values indicate better sampling quality.
@@ E.4 Further Discussions on Task-specific Schedule Preferences
Among the six tasks we experimented on, BBH [28] is only one that focuses on general reasoning problems. As for GSM8K [10], although it is mathematics-focused, its answers use detailed natural language explanations compared to other math benchmarks. This might be the reason why sustained refinement in the middle phases are preferred. Considering this similarity, it is reasonable that GSM8K [10] and BBH [28] have similar preference on schedules although it is difficult to mathematically deduce the exact expression of the energy functionals.
@@ E.5 Additional Samples
Table 8: Comparison of beta reparameterized and manually designed schedules on math-reasoning task Hendrycks Math [11]. Sampling steps set as 64.
@@
Table 9: Comparison of beta reparameterized and manually designed schedules on code-generating task HumanEval [9]. Sampling steps set as 64.

### v2

@@ Appendix B An Intuitive Explanation of Geodesics
This appendix provides an intuitive introduction to geodesics and exponential maps to clarify the geometric interpretation of MDM in [jo2025]. Readers seeking formal mathematical definitions of these differential geometry concepts may refer to [ay2017information] for complete technical specifications.
@@
Recent theoretical advances [jo2025] reveal that MDM’s conditional probability flow in Eq. (3) forms exactly the geodesic curve in spherical geometry (see Fig. 6). The interpolation schedule $\gamma_{t}$ is uniquely determined by the mask schedule $\alpha_{t}$ as $\gamma_{t}=\frac{2}{\pi}\arcsin\sqrt{\alpha_{t}}$, which is equivalent with Condition 3.4.
@@ C.2 Proof of the Invariance of Training Loss to the Mask Schedule
Different equivalence expressions of the training loss of MDM has been proved invariant to the choice of $\alpha_{t}$ in multiple works [kingma2021variational, shi2024simplified, mdlm]. We adapt a proof from [mdlm] by examining the negative evidence lower bound (NELBO) through token-level denoising components:
@@
Proposition C.2 (Schedule Invariance, Proof Adapted from [mdlm]).
@@ C.4 Decomposition of Conditional Kinetic Energy along Sequence Dimension
The pivotal decomposition in $(1)$ leverages two structural properties: (i) The conditional probability flow (Eq. (3)) exhibits token-wise independence, and (ii) The conditional rate matrix (Lemma C.3) nullifies transitions altering multiple tokens simultaneously. These enable reduction of full-sequence transitions to single-token operations, with remaining n1n-1 tokens contributing constant combinatorial factors. This structural property persists across various DPF implementations including MDM and Discrete Flow Matching [DiscreteFlowMatching, DiscretePaths], validating the conditional kinetic energy as theoretically sound surrogate objective. Notably, standard kinetic energy $\mathcal{E}$ lacks such decomposition due to $p_{t}(z)$’s dependence on cross-token correlations.
@@ Appendix D Proof of Main Results
Our argument leverages a foundational decomposition from [radd] regarding concrete score representations:
@@
where $(1)$ leverages orthogonality $p_{0}\cdot p_{1}=0$. Therefore, we proved that MDM with schedule schedule $\alpha_{t}^{\star}$ generates the minimal-length curve as well as minimal-energy conditional probability path, validating and generalizing the conclusion in [jo2025] from a energy perspective.
@@ D.5 Optimality in the Discretized Case
where $Q_{t|1}(\alpha_{t}^{\star})$ refer to conditional rate matrix in MDM case derived in Appendix C.3 under Condition 3.4. We adapt the following key result from prior analysis  [DiscretePaths]of the $\gamma_{t}=t$ case:
@@ E.2 Standard Benchmarks and Evaluation Settings
Building upon established practices in LLM evaluation [chu2024qwen2, qwen2.5, llada], we evaluate performance across key dimensions including: general ability (BBH [metric:bbh]), mathematics (GSM8K [metric:gsm8k], Hendrycks MATH [metric:math], Minerva MATH [lewkowycz2022solving]), and code generation (MBPP [metric:mbpp], HumanEval [metric:humaneval]). Evaluation follows the conditional generation paradigm, where models produce completions given task prompts, with performance quantified through exact match or other domain-specific evaluation metrics.
Our implementation leverages the open-source pretrained weights and evaluation toolkit from LLaDA [llada], modifying only the mask schedule that governs the iterative unmasking process. The mask schedule affects the number of tokens unmasked at each step, with certain schedules permitting zero-token unmasking during the process. Therefore, generation quality discrepancies occur even when sampling steps are set as the sequence length. All experiments can be efficiently conducted on a single NVIDIA A800 GPU.
@@ E.3 Additional Results and Raw Data
Fig. 7 shows the result of our main experiments on benchmark BBH [metric:bbh] and GSM8K [metric:gsm8k], where our beta-parameterized schedules exhibit comparable yet not better performance than the linear schedule.
@@
Figure 7: Performance evaluation of energy-optimized schedules on BBH [metric:bbh] and GSM8K [metric:gsm8k] where our beta-parameterized schedules exhibit comparable yet not better performance than the linear schedule. Each panel corresponds to a distinct benchmark. The x-axis displays sampling steps on a logarithmic scale, while the y-axis quantifies task performance, where higher values denote superior generation quality.
Table 2: Performance evaluation of beta-parameterized schedules on MBPP [metric:mbpp] benchmark. All experiments fix generation length at 256 and higher values indicate better sampling quality.
@@
Table 3: Performance evaluation of beta-parameterized schedules on HumanEval [metric:humaneval] benchmark. All experiments fix generation length at 256 and higher values indicate better sampling quality.
@@
Table 4: Performance evaluation of beta-parameterized schedules on BBH [metric:bbh] benchmark. All experiments fix generation length at 128 and higher values indicate better sampling quality.
@@
Table 5: Performance evaluation of beta-parameterized schedules on GSM8K [metric:gsm8k] benchmark. All experiments fix generation length at 128 and higher values indicate better sampling quality.
@@
Table 6: Performance evaluation of beta-parameterized schedules on Hendrycks Math [metric:math] benchmark. All experiments fix generation length at 256 and higher values indicate better sampling quality.
@@
Table 7: Performance evaluation of beta-parameterized schedules on Minerva Math [lewkowycz2022solving] benchmark. All experiments fix generation length at 256 and higher values indicate better sampling quality.
@@ E.4 Further Discussions on Task-specific Schedule Preferences
Among the six tasks we experimented on, BBH [metric:bbh] is only one that focuses on general reasoning problems. As for GSM8K [metric:gsm8k], although it is mathematics-focused, its answers use detailed natural language explanations compared to other math benchmarks. This might be the reason why sustained refinement in the middle phases are preferred. Considering this similarity, it is reasonable that GSM8K [metric:gsm8k] and BBH [metric:bbh] have similar preference on schedules although it is difficult to mathematically deduce the exact expression of the energy functionals.
@@ E.5 Additional Samples
Table 8: Comparison of beta reparameterized and manually designed schedules on math-reasoning task Hendrycks Math [metric:math]. Sampling steps set as 64.
@@
Table 9: Comparison of beta reparameterized and manually designed schedules on code-generating task HumanEval [metric:humaneval]. Sampling steps set as 64.